### v1

# PDMLP: Patch-based Decomposed MLP for Long-Term Time Series Forecasting

Abstract
Recent studies have attempted to refine the Transformer architecture to demonstrate its effectiveness in Long-Term Time Series Forecasting (LTSF) tasks. Despite surpassing many linear forecasting models with ever-improving performance, we remain skeptical of Transformers as a solution for LTSF. We attribute the effectiveness of these models largely to the adopted Patch mechanism, which enhances sequence locality to an extent yet fails to fully address the loss of temporal information inherent to the permutation-invariant self-attention mechanism. Further investigation suggests that simple linear layers augmented with the Patch mechanism may outperform complex Transformer-based LTSF models. Moreover, diverging from models that use channel independence, our research underscores the importance of cross-variable interactions in enhancing the performance of multivariate time series forecasting. The interaction information between variables is highly valuable but has been misapplied in past studies, leading to suboptimal cross-variable models. Based on these insights, we propose a novel and simple Patch-based Decomposed MLP (PDMLP) for LTSF tasks. Specifically, we employ simple moving averages to extract smooth components and noise-containing residuals from time series data, engaging in semantic information interchange through channel mixing and specializing in random noise with channel independence processing. The PDMLP model consistently achieves state-of-the-art results on several real-world datasets. We hope this surprising finding will spur new research directions in the LTSF field and pave the way for more efficient and concise solutions.
1 Introduction
Figure 1: The self-attention scores of a 2-layer Transformer with different Patch size trained on ETTh1. We follow the setup of PathcTST Nie et al. (2022), retaining only the Encoder while replacing the Decoder with a simple MLP, and using a channel independent approach. A patch size of 1 is equivalent to the original Transformer, indicating that time series data often exhibits a trend of being segmented into patches Zhang & Yan (2022); Tang & Zhang (2023), and an increase in patch size can mitigate this to some extent.
Long-term Time Series Forecasting (LTSF) is a critical area of research in the fields of statistics and machine learning, aimed at using historical data to predict the future of one or more variables over a certain period Oreshkin et al. (2019); Gong et al. (2023); Lin et al. (2023). Time series data are organized in chronological order and reveal underlying dynamic patterns, be they cyclical or non-cyclical Cryer (1986). Such forecasts play a vital role in various sectors, including biomedicine Liu et al. (2018), economics and finance Patton (2013), electricity Zhou et al. (2021), and transportation Yin & Shang (2016).
Multivariate Time Series (MTS) consist of multiple variables recorded at the same time point, where each dimension may represent an individual univariate time series or be considered as a signal with multiple channels. With the advancement of deep learning, a number of models have been developed to enhance the performance of MTS forecasting Borovykh et al. (2017); Bai et al. (2018); Liu et al. (2020). In particular, recent models based on the Transformer Vaswani et al. (2017) architecture have demonstrated significant potential in capturing long-term dependencies Radford et al. (2018); Devlin et al. (2018). In recent years, Transformers have emerged as a leading architecture in time series forecasting Lim et al. (2021); Liu et al. (2022); Shen et al. (2023), initially applied in the field of Natural Language Processing (NLP) Radford et al. (2019); Brown et al. (2020) and subsequently expanded to other domains such as Computer Vision (CV) Liu et al. (2021b); Feichtenhofer et al. (2022) through the adaptation of the Patch method, becoming a versatile framework.
Originally, these models utilized a channel mixing approach Li et al. (2023b), which involves projecting vectors from different channels recorded at the same time point into an embedding space and blending the information Zhou et al. (2021); Tang & Zhang (2023) ,and many models adopt the idea of time series decomposition Wu et al. (2021); Zhou et al. (2022); Wang et al. (2023). Nonetheless, recent studies have shown that a channel independence model might be more effective Nie et al. (2022); Han et al. (2023). Channel independence means that each input token contains information from only one channel, and intuitively, the models treat the MTS as separate univariate series for individual processing. These studies suggest that for LTSF tasks, an emphasis on channel independence may be more effective than channel mixing strategies Gong et al. (2023).
Moreover, the effectiveness of Transformers in LTSF tasks has been challenged by a reconsideration of the Multi-Layer Perceptron (MLP) Zeng et al. (2023); Ekambaram et al. (2023); Chen et al. (2023), whose surprisingly simple architecture has outperformed all previous Transformer models. This raises a compelling question: Are Transformers effective for LTSF tasks? In response to this skepticism, recent Transformer-based models have adopted patch-based representations and achieved noteworthy performance in LTSF Nie et al. (2022); Chen et al. (2024).
This study raises three critical inquiries:
  • Is the method of channel mixing truly ineffective in MTS forecasting?
  • Can simply decomposing the original time series truly better predict the trend and seasonal components?
  • Does the impressive performance of the Patch-based Transformer derive from the inherent strength of the Transformer architecture, or is it merely due to the use of patches as the input representation?
Figure 2: Experimental results of Patch Transformer on the ETTh2 dataset. (a) Maintain all other parameters constant, and present the MSE outcomes for four forecast lengths with only the input length altered. (b) Keeping all other parameters constant and only altering the size of the patch, the MSE results for a forecast length of 720 with five different input lengths. (c) Maintaining all other parameters unchanged and only varying the patch size, the MSE results for five different $d_{model}$ values with both input and forecast lengths set to 720.
In this paper, we specifically discuss the three issues mentioned above and introduce the PDMLP, a novel and concise Patch-based decomposed MLP model tailored for LTSF task. PDMLP is entirely based on fully connected networks and incorporates the concept of patche. Specifically, we employ a patch methodology to embed the original sequences into representational spaces, then extract the smooth components and noisy residuals of the series using a simple moving average technique for separate processing. We independently process stochastic noise across channels and facilitate semantic information interchange between variables through channel mixing. Upon evaluation on numerous real-world datasets, the PDMLP model demonstrated state-of-the-art (SOTA) performance. To be precise, our contributions can be summarized in three aspects:
  • We analyze the effectiveness of Patches in time series forecasting and propose a Multi-Scale Patch Embedding (MPE) approach. Unlike previous embedding methods that used a single linear layer, MPE is capable of capturing multiscale relationships between input sequences.
  • We introduce a novel entirely MLP-based model, named PDMLP. By utilizing moving averages, it performs the decomposition of latent vector and adopts a different approach to channel mixing for semantic information exchange across variables.
  • Our experiments across a wide range of datasets in various fields show that PDMLP consistently achieves SOTA performance across multiple forecasting benchmarks. Moreover, we conduct an extensive analysis of Patch-based methods with the aim of charting new directions for future research in time series forecasting.
2 Patch for Long-Term Time Series Forecasting
Why is Patch effective in time series forecasting? Lee et al. (2023); Zhong et al. (2023) We have studied a patch-based Transformer model adopting channel independent mechanisms, and we set the decoder layer as a simple single-layer MLP for time series forecastine Nie et al. (2022). As shown in the Figure  2(b), we can clearly see that for the same input length, as the size of the patch increases, the overall Mean Squared Error (MSE) of the model exhibits a trend of decreasing and then increasing, or decreasing and then stabilizing. As the input length increases, the patch size required for the model to achieve optimal performance also gradually increases. In the original input data, the effect of Attention is so poor that it cannot even outperform a simple single-layer MLP, which raises the question of whether Attention might not be the optimal choice for time series modeling to some extent, as it cannot handle long-term time series well.
Compared to textual data in the NLP field, original time series data contains many redundant features due to high-frequency sampling, and is easily influenced by noise to some extent during the sampling process Tang & Zhang (2022). We believe that for data with excessive noise in the original data, the original Transformer’s Attention mechanism is not the optimal choice as it cannot effectively eliminate the noise. Sparse Attention performs better because this sparse mechanism can effectively mitigate the impact of noise to some extent Li et al. (2019); Wu et al. (2020); Liu et al. (2021a), but it may also attenuate the impact of original features.
Patch, on the other hand, compresses the data, reduces the dimensionality of the input data, and decreases redundant features. Additionally, patch provides a certain degree of smoothing, which can reduce the influence of outliers to some extent, and help filter out fluctuations and random noise, retaining more stable and representative information. In the CV field, patch also works well in ViT Dosovitskiy et al. (2020), MAE He et al. (2022). Time series data usually contains patterns at different scales, and patch provides a modeling of short-term time series, enhancing local information in the sequence, allowing the model to better learn and capture local features. Therefore, we believe that the effectiveness of Transformer is not due to the effect of Attention, but rather due to the presence of patch.
Figure 3: Overall structure of PDMLP. (a) Raw time series of different variates are independently embedded as tokens. (b)The Feature Decomposition utilizing moving averages to decompose the tokens into the smooth components and noisy residuals. (c) The MLP processed sequences in two ways: intra-variable and inter-variable. (d) A projection maps the latent vectors back to future sequences.
So, is a larger patch always better? In Figure  2(b), a larger patch does not seem to achieve better results, which may be partly influenced by the size of the hidden layer ($d_{model}$). With a larger patch size, more time points are projected into a fixed-size $d_{model}$, which may lead to excessive compression. As shown in Figure  2(c), as $d_{model}$ increases, the performance of the model gradually improves, indicating that larger patches perform better. In the extreme case, a patch that treats the entire input sequence as one patch and projects it into a vector is similar to iTransformer Liu et al. (2023). However, a larger $d_{model}$ indicates that there are more parameters for the model to learn, which can easily lead to underfitting and result in decreased model performance.
3 PDMLP
The problem of multivariate time series forecasting is to input the historical observations $\mathcal{X}=\left\{x_{1},\cdots,x_{L}|x_{i}\in\mathbb{R}^{M}\right\}$, and the output is to predict corresponding future sequence $\mathcal{X}=\left\{x_{L+1},\cdots,x_{L+T}|x_{i}\in\mathbb{R}^{M}\right\}$, where $L$ and $T$ are the lengths of input and output sequences respectively, and $M$ is the dimension of variates.
As shown in Figure  3, PDMLP consists of four network components: Multi-Scale Patch Embedding layer, Feature Decomposition layer, Multi-Layer Perceptron (MLP) layer, Projection layer. The Multi-Scale Patch Embedding layer embeds the multivariate time series into latent space. The Feature Decomposition layer decomposes the latent vector into the smooth components and noisy residuals, then operate separately through MLP layer. Finally, the latent vector are mapped back to the feature space through the Projection layer to obtain the future sequences $\hat{\mathcal{X}}$. Next, we will introduce the above modules separately.
3.1 Multi-Scale Patch Embedding
Time series analysis relies on the accurate identification of local information within the sequence and optimization of the model. The use of patches can provide the model with a local view of the time series in the short term, thereby enhancing the representation of local information within the sequence and enabling the model to learn and capture these local features more precisely. However, traditional methods typically employ patches of a single-scale to embed the raw time series, which makes the model inclined to learn single-scale temporal relationships while neglecting the multi-scale nature and complexity of time series data.
In practical applications, the single-scale patch strategy may lead to the model capturing inaccurate or incomplete local features. This is because the data characteristics of different time series often have variation, and a patch of the same scale cannot universally adapt to all types of time series. For example, a time series containing multiple cyclical patterns, a single-scale patch cannot effectively identify and learn the cyclical features present at different frequencies.
To capture local information in more detail and to fully understand the temporal relationships within the time series, we adopt multi-scale patche. This includes shorter patch to capture local high-frequency patterns, as well as longer patch to unearth long-term seasonal, trends and periodic fluctuations. Through the introduction of multi-scale patch, the model can flexibly learn representative features over different lengths of time spans, thus enhancing predictive accuracy and model generalization capabilities.
To decompose the multivariate time series $\mathcal{X}$ into a univariate series $x$, we have developed a suite of patches $\mathcal{P}$ across various scales to process $x$. For a particular scale $p\in\mathcal{P}$, $x$ is first divided into non-overlapping patches where the length of patch corresponds to $p$, thus the patching process yields the sequence of patches $x_{p}\in\mathbb{R}^{N\times p}$, where $N$ is the number of patches. Subsequently, we employ a single-layer linear layer to embed the patches $x_{p}$, resulting in latent vectors $x_{e}\in\mathbb{R}^{N\times d}$, $d$ is the embedding dimension, the $d$ corresponding to different scales patches can be different. These latent vectors are then unfolded and concatenated with latent vectors obtained from patches at alternative scales, thus assembling the ultimate embedding vector $X\in\mathbb{R}^{1\times d_{model}}$, $d_{model}$ is the final embedding dimension input into the model. This multi-scale patch strategy permits the model to capture and learn the intrinsic dynamics of time series at different levels, thereby providing more precise insights when forecasting future trends and patterns.
3.2 Feature Decomposition
Time series often contain complex temporal patterns, including seasonal fluctuations, trends, and other irregular influencing factors. Many studies attempt to decompose time series into these fundamental components through decomposition methods, to achieve better understanding and forecasting. Although decomposition is a powerful tool, traditional methods may encounter some difficulties when dealing with the raw sequences, especially when there are complex and mixed patterns in the sequences. These methods often struggle to accurately separate clear trends and seasonal components, particularly when the data contains substantial noise or nonlinear components.
In response to this issue, we propose a different idea, instead of directly decomposing the original sequence, we decompose the sequence’s embedding vector. Embedding vector are representations formed by mapping time series to a high-dimensional space, and these representations often capture the core information of the original data. By decomposing embedding vector, we can distinguish smoother components and noise-containing residuals, which helps eliminate the interference of random fluctuations on analysis and forecasting. In practice, we use the operation of Average Pooling ($AvgPool$) to smooth the time series, aiming to reduce random fluctuations and noise in the data. Moreover, in order to maintain the length of the time series unchanged while smoothing, we applied a padding operation.
$\begin{split}X_{s}&=AvgPool(X)\\ X_{r}&=X-X_{s}\end{split}$ (1)
$X_{s}$ and $X_{r}$ respectively represent the extracted smooth components and residual components. By operating on embedding vector rather than the original sequence, the model can extract and identify the fundamental components of time series more precisely, enabling a better understanding of the intrinsic structure of time series, and ultimately, improving the accuracy of time series forecasting.
Figure 4: Overall structure of MLP layer. (a) The embedded vectors first interact with the temporal information within the variable through the Intra-Variable MLP. (b) Then interact with the feature domain information between variables through the Intra-Variable MLP. (c) Subsequently, they are multiplied by the input of the Inter-Variable MLP using a dot-product approach. (d) Finally, they are added to the initial input of the MLP Layer using skip connections.
3.3 MLP layer
In the context of MTS forecasting, the MLP layer alternately applies MLPs within Intra-Variable (time domain) and Inter-Variable(feature domain) to enhance predictive performance. The specific architecture, as depicted in Figure  4, can be described in detail as follows:
Intra-Variable MLP: This component is focused on the identification of time-correlated patterns within time series. Specifically, it utilizes a network architecture composed of fully connected layers, nonlinear activation functions, and dropout. Parameters are applied in the temporal domain and are shared across inter-variable. Simple linear models are easy to understand and implement, and have been proven to learn complex time patterns and capture more complex time dependencies.
Inter-Variable MLP: Complementary to the intra-variable MLP, the inter-variable MLP aims to model the mutual influences among MTS variables. This component also consists of fully connected layers, activation functions, and dropout, applying fully connected layers in the feature domain to achieve parameter sharing within the intra-variable. To enhance the cross-variable interactivity, we utilize a dot product mechanism, integrating the dot product results between the MLP outputs and inputs, which enhances the model’s nonlinear representational capability.
Residual Connections: Residual connections He et al. (2016) are applied after each MLP, enabling the model to learn deeper architectures more efficiently. These connections not only help to mitigate the problems of vanishing gradients in deep networks but also provide the model with a shortcut path that ensures key temporal or feature information is not overlooked.
While the minimalist architecture of PDMLP may not appear as eye-catching as some of the recently proposed, more complex Transformer models, empirical results indicate that PDMLP exhibits competitive performance, both in terms of training efficiency and accuracy on standard benchmark tests, relative to SOTA models.
Table 1: Full results of the Long-Term Time Series Forecasting task. For all baselines, we adhere to the setting of the iTransformer with an input sequence length of 96. We conducted comparisons among an array of competitive models during various forecast horizons, and the outcomes were obtained from the iTransformer. Avg denotes the average result across all four prediction lengths. We denote the optimal results with boldface, and the second-best with an underline for clarity.
Methods PDMLP iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear SCINet FEDformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.316 0.357 0.334 0.368 0.355 0.376 0.329 0.367 0.404 0.426 0.364 0.387 0.338 0.375 0.345 0.372 0.418 0.438 0.379 0.419
192 0.357 0.378 0.377 0.391 0.391 0.392 0.367 0.385 0.450 0.451 0.398 0.404 0.374 0.387 0.380 0.389 0.439 0.450 0.426 0.441
336 0.396 0.402 0.426 0.420 0.424 0.415 0.399 0.410 0.532 0.515 0.428 0.425 0.410 0.411 0.413 0.413 0.490 0.485 0.445 0.459
720 0.462 0.439 0.491 0.459 0.487 0.450 0.454 0.439 0.666 0.589 0.487 0.461 0.478 0.450 0.474 0.453 0.595 0.550 0.543 0.490
Avg 0.382 0.394 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.485 0.481 0.448 0.452
ETTm2 96 0.175 0.258 0.180 0.264 0.182 0.265 0.175 0.259 0.287 0.366 0.207 0.305 0.187 0.267 0.193 0.292 0.286 0.377 0.203 0.287
192 0.240 0.302 0.250 0.309 0.246 0.304 0.241 0.302 0.414 0.492 0.290 0.364 0.249 0.309 0.284 0.362 0.399 0.445 0.269 0.328
336 0.306 0.344 0.311 0.348 0.307 0.342 0.305 0.343 0.597 0.542 0.377 0.422 0.321 0.351 0.369 0.427 0.637 0.591 0.325 0.366
720 0.404 0.403 0.412 0.407 0.407 0.398 0.402 0.400 1.730 1.042 0.558 0.524 0.408 0.403 0.554 0.522 0.960 0.735 0.421 0.415
Avg 0.281 0.326 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.571 0.537 0.305 0.349
ETTh1 96 0.382 0.399 0.386 0.405 0.386 0.395 0.414 0.419 0.423 0.448 0.479 0.464 0.384 0.402 0.386 0.400 0.654 0.599 0.376 0.419
192 0.437 0.431 0.441 0.436 0.437 0.424 0.460 0.445 0.471 0.474 0.525 0.492 0.436 0.429 0.437 0.432 0.719 0.631 0.420 0.448
336 0.475 0.450 0.487 0.458 0.479 0.446 0.501 0.466 0.570 0.546 0.565 0.515 0.491 0.469 0.481 0.459 0.778 0.659 0.459 0.465
720 0.492 0.474 0.503 0.491 0.481 0.470 0.500 0.488 0.653 0.621 0.594 0.558 0.521 0.500 0.519 0.516 0.836 0.699 0.506 0.507
Avg 0.446 0.438 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.552 0.541 0.507 0.458 0.450 0.456 0.452 0.747 0.647 0.440 0.460
ETTh2 96 0.288 0.340 0.297 0.349 0.288 0.338 0.302 0.348 0.745 0.584 0.400 0.440 0.340 0.374 0.333 0.387 0.707 0.621 0.358 0.397
192 0.371 0.396 0.380 0.400 0.374 0.390 0.388 0.400 0.877 0.656 0.528 0.509 0.402 0.414 0.477 0.476 0.860 0.689 0.429 0.439
336 0.412 0.429 0.428 0.432 0.415 0.426 0.426 0.433 1.043 0.731 0.643 0.571 0.452 0.452 0.594 0.541 1.000 0.744 0.496 0.487
720 0.421 0.441 0.427 0.445 0.420 0.440 0.431 0.446 1.104 0.763 0.874 0.679 0.462 0.468 0.831 0.657 1.249 0.838 0.463 0.474
Avg 0.373 0.401 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.954 0.723 0.437 0.449
ECL 96 0.147 0.246 0.148 0.240 0.201 0.281 0.195 0.285 0.219 0.314 0.237 0.329 0.168 0.272 0.197 0.282 0.247 0.345 0.193 0.308
192 0.166 0.261 0.162 0.253 0.201 0.283 0.199 0.289 0.231 0.322 0.236 0.330 0.184 0.289 0.196 0.285 0.257 0.355 0.201 0.315
336 0.184 0.279 0.178 0.269 0.215 0.298 0.215 0.305 0.246 0.337 0.249 0.344 0.198 0.300 0.209 0.301 0.269 0.369 0.214 0.329
720 0.222 0.311 0.225 0.317 0.257 0.331 0.256 0.337 0.280 0.363 0.284 0.373 0.220 0.320 0.245 0.333 0.299 0.390 0.246 0.355
Avg 0.179 0.274 0.178 0.270 0.219 0.298 0.216 0.304 0.244 0.334 0.251 0.344 0.192 0.295 0.212 0.300 0.268 0.365 0.214 0.327
Exchange 96 0.085 0.204 0.086 0.206 0.093 0.217 0.088 0.205 0.256 0.367 0.094 0.218 0.107 0.234 0.088 0.218 0.267 0.396 0.148 0.278
192 0.174 0.298 0.177 0.299 0.184 0.307 0.176 0.299 0.470 0.509 0.184 0.307 0.226 0.344 0.176 0.315 0.351 0.459 0.271 0.315
336 0.317 0.411 0.331 0.417 0.351 0.432 0.301 0.397 1.268 0.883 0.349 0.431 0.367 0.448 0.313 0.427 1.324 0.853 0.460 0.427
720 0.820 0.681 0.847 0.691 0.886 0.714 0.901 0.714 1.767 1.068 0.852 0.698 0.964 0.746 0.839 0.695 1.058 0.797 1.195 0.695
Avg 0.349 0.398 0.360 0.403 0.378 0.417 0.367 0.404 0.940 0.707 0.370 0.413 0.416 0.443 0.354 0.414 0.750 0.626 0.519 0.429
Traffic 96 0.424 0.278 0.395 0.268 0.649 0.389 0.544 0.359 0.522 0.290 0.805 0.493 0.593 0.321 0.650 0.396 0.788 0.499 0.587 0.366
192 0.441 0.289 0.417 0.276 0.601 0.366 0.540 0.354 0.530 0.293 0.756 0.474 0.617 0.336 0.598 0.370 0.789 0.505 0.604 0.373
336 0.461 0.293 0.443 0.283 0.609 0.369 0.551 0.358 0.558 0.305 0.762 0.477 0.629 0.336 0.605 0.373 0.797 0.508 0.621 0.283
720 0.482 0.307 0.467 0.302 0.647 0.387 0.586 0.375 0.589 0.328 0.719 0.449 0.640 0.350 0.645 0.394 0.841 0.523 0.626 0.382
Avg 0.452 0.291 0.428 0.282 0.626 0.378 0.555 0.362 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383 0.804 0.509 0.610 0.376
Weather 96 0.157 0.209 0.174 0.214 0.192 0.232 0.177 0.218 0.158 0.230 0.202 0.261 0.172 0.220 0.196 0.255 0.221 0.306 0.217 0.296
192 0.206 0.251 0.221 0.254 0.240 0.271 0.225 0.259 0.206 0.277 0.242 0.298 0.219 0.261 0.237 0.296 0.261 0.340 0.276 0.336
336 0.267 0.293 0.278 0.296 0.292 0.307 0.278 0.297 0.272 0.335 0.287 0.335 0.280 0.306 0.283 0.335 0.309 0.378 0.339 0.380
720 0.347 0.346 0.358 0.349 0.364 0.353 0.354 0.348 0.398 0.418 0.351 0.386 0.365 0.359 0.345 0.381 0.377 0.427 0.403 0.428
Avg 0.244 0.274 0.258 0.279 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.292 0.363 0.309 0.360
Solar-Energy 96 0.202 0.235 0.203 0.237 0.322 0.339 0.234 0.286 0.310 0.331 0.312 0.399 0.250 0.292 0.290 0.378 0.237 0.344 0.242 0.342
192 0.239 0.262 0.233 0.261 0.359 0.356 0.267 0.310 0.734 0.725 0.339 0.416 0.296 0.318 0.320 0.398 0.280 0.380 0.254 0.272
336 0.251 0.272 0.248 0.273 0.397 0.369 0.290 0.315 0.750 0.735 0.368 0.430 0.319 0.330 0.353 0.415 0.304 0.389 0.282 0.376
720 0.248 0.275 0.249 0.275 0.397 0.356 0.289 0.317 0.769 0.765 0.370 0.425 0.338 0.337 0.356 0.413 0.308 0.388 0.357 0.427
Avg 0.235 0.261 0.233 0.262 0.369 0.270 0.307 0.641 0.6398 0.347 0.417 0.301 0.319 0.330 0.401 0.282 0.375 0.281 0.291 0.381
$\mathrm{1^{st}Count}$ 22 20 11 11 3 11 5 3 0 0 0 0 0 0 0 0 0 0 4 0
Table 2: Ablations of PDMLP. We replace the components of PDMLP one by one and explore the performance of different MLPs. $Avg$ represents the average results across all four forecast lengths. For clarity, the best results are highlighted in bold.
Dateset ECL Solar-Energy Traffic
Model Metric 96 192 336 720 Avg 96 192 336 720 Avg 96 192 336 720 Avg
PDMLP MSE 0.147 0.166 0.184 0.222 0.179 0.202 0.239 0.251 0.248 0.235 0.424 0.441 0.461 0.482 0.452
MAE 0.246 0.261 0.279 0.311 0.274 0.235 0.262 0.272 0.275 0.261 0.278 0.289 0.293 0.307 0.291
PDMLP1 MSE 0.152 0.168 0.192 0.237 0.187 0.217 0.252 0.281 0.274 0.256 0.469 0.484 0.508 0.548 0.502
MAE 0.252 0.265 0.291 0.325 0.283 0.260 0.282 0.300 0.291 0.283 0.324 0.332 0.354 0.365 0.343
PDMLP2 MSE 0.151 0.169 0.187 0.226 0.183 0.210 0.243 0.263 0.261 0.244 0.522 0.525 0.539 0.570 0.539
MAE 0.254 0.268 0.285 0.319 0.281 0.256 0.276 0.290 0.288 0.277 0.342 0.353 0.360 0.451 0.376
PDMLP3 MSE 0.174 0.180 0.201 0.242 0.199 0.222 0.253 0.291 0.289 0.263 0.448 0.459 0.491 0.523 0.480
MAE 0.257 0.263 0.283 0.317 0.280 0.258 0.277 0.300 0.297 0.283 0.290 0.292 0.309 0.327 0.304
DPMLP MSE 0.151 0.168 0.189 0.227 0.183 0.218 0.249 0.263 0.271 0.250 0.480 0.495 0.514 0.554 0.510
MAE 0.251 0.265 0.289 0.320 0.281 0.270 0.292 0.304 0.299 0.291 0.328 0.339 0.346 0.362 0.343
PMLP MSE 0.153 0.170 0.194 0.242 0.189 0.219 0.257 0.274 0.267 0.254 0.474 0.482 0.512 0.557 0.506
MAE 0.254 0.266 0.292 0.330 0.285 0.249 0.288 0.295 0.289 0.280 0.328 0.329 0.342 0.369 0.342
PMLP1 MSE 0.153 0.170 0.197 0.245 0.191 0.225 0.268 0.287 0.285 0.266 0.484 0.491 0.516 0.554 0.510
MAE 0.252 0.268 0.296 0.335 0.287 0.262 0.290 0.303 0.301 0.289 0.334 0.336 0.350 0.370 0.347
PMLP2 MSE 0.153 0.169 0.188 0.227 0.184 0.221 0.266 0.267 0.269 0.255 0.526 0.531 0.528 0.556 0.535
MAE 0.257 0.268 0.288 0.319 0.283 0.248 0.290 0.280 0.283 0.275 0.345 0.350 0.349 0.363 0.351
PMLP3 MSE 0.179 0.185 0.201 0.243 0.202 0.234 0.255 0.290 0.328 0.276 0.455 0.471 0.490 0.517 0.483
MAE 0.260 0.267 0.284 0.317 0.282 0.269 0.278 0.299 0.321 0.291 0.293 0.298 0.308 0.322 0.305
3.4 Projection and Loss Function
We adopt a channel independent approach to implement $\mathrm{Projection}:\mathbb{R}^{d_{model}}\to\mathbb{R}^{T}$ by MLP, where each channel shares parameters.
Our loss function is calculated by the Mean Square Error (MSE) between the model prediction $\hat{x}$ and the ground truth $x$:
$\mathcal{L}=\frac{1}{M}\sum_{M}^{i=1}\left\|\hat{x}^{i}_{L+1:L+T}-x^{i}_{L+1:L+T}\right\|^{2}_{2}$ (2)
and the loss is propagated back from the Projection output across the entire model.
4 Experiments
We comprehensively evaluated the proposed PDMLP in various time series forecasting applications, verifying the generality of the proposed framework, and have further delved into the investigation of the effectiveness of individual components of PDMLP when applied in LSTF tasks.
Datasets: We evaluated the performance of our proposed PDMLP on 8 commonly used LSTF benchmark datasets: Solar Energy Lai et al. (2018), Weather, Traffic, Electricity (ECL) Wu et al. (2021), and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) Zhou et al. (2021).
Baselines and Metrics: We selected nine widely acknowledged SOTA forecasting models for our benchmarking analysis, which comprises the Transformer-based models: iTransformer Liu et al. (2023), PatchTST Nie et al. (2022), Crossformer Zhang & Yan (2022), and FEDformer Zhou et al. (2022); the CNN-based models: TimeNet Wu et al. (2022); and the significant MLP-based models: DLinear Zeng et al. (2023), TiDE Das et al. (2023), and RLinear Li et al. (2023a). To evaluate the performance of these models, we used widely used evaluation metrics: MSE and Mean Absolute Error (MAE).
4.1 Multivariate Long-Term Forecasting
Table  1 presents the results of MTS forecasting, with lower MSE/MAE values indicating greater predictive accuracy. Overall, the model we propose has achieved a surprisingly efficacious outcome, realizing optimal performance across all or the majority of forecast horizons. Out of a total of 90 metrics, a remarkable count of 42 SOTA results were attained, surpassing the second-ranked iTransformer by nearly 50%. This suggests that the Attention mechanism may not necessarily be the optimal choice for LSTF tasks, as even simple linear models are capable of achieving impressive results. Further analysis indicates that channel independence models, represented by PatchTST, often failed to perform to their potential, thus implying the indispensable role of inter-variable interactions in MTS analysis. Notably, despite explicitly capturing multivariate correlations, the performance of Crossformer did not meet expectations, underscoring that inappropriate utilization of inter-variable correlations can negatively impact model efficacy. Hence, we conclude that skillful leverage of interactions between variables is vital and, in certain scenarios, simple linear models may suffice to deliver outstanding performance compared to the more complex Transformer architectures.
4.2 Model Analysis
4.2.1 Ablation Study
As shown in the Table  2, we investigate the impact of each module of PDMLP on performance. DPMLP follows the conventional decomposition approach, which involves decomposing the time series first and then embedding it before separately handling the prediction of the trend and seasonal components. PMLP eliminates the decomposition module, directly inputs the embedded original sequence into the model for prediction. In both PDMLP1 and PMLP1, the MPE is removed, and the embedding method is altered to a single linear layer. PDMLP2 and PMLP2, on the other hand, eliminate the dot product in the MLP among variables and replace it with a simple residual connection, while PDMLP3 and PMLP3 completely remove the MLP among variables, canceling the interaction between them. We only change the corresponding modules and keep all other settings unchanged, with the input length still set to 96.
It can be observed that conventional decomposition methods do not perform well, which is attributable to the excessively complex temporal relationships inherent in the original time series that cannot be effectively deconstructed through simple decomposition techniques. In contrast, latent vector decomposition can adeptly circumvent this issue. Furthermore, the observed performance deterioration upon the removal of the MPE indicates that MPE plays a significant role in learning the various temporal relationships. Additionally, it is noteworthy that the dot product method outperforms simply because it enhances the interactiveness between variables, while mere addition does not confer this advantage. Correspondingly, models that lack variable interactiveness experience a predictable decline in performance, which reiterates the critical importance of appropriately leveraging the interrelationships between variables.
4.2.2 Increasing look-back Window
Figure 5: Forecasting performance (MSE) with varying look-back windows on 3 datasets: ETTh1, ETTm2, and Weather. The look-back windows are selected to be $L=\{192,288,384,480,576,672,768\}$, and the prediction horizons are $T=\{192,720\}$.
In principle, a longer input sequence might enhance model performance, primarily due to the provision of a richer historical context that assists the model in more effectively learning and identifying long-term patterns within time series. Generally, a robust LSTF model, equipped with substantial capability to extract temporal relationships, is expected to yield improved outcomes with extended input lengths. However, longer inputs necessitate the model’s heightened ability to capture long-term dependencies, as a deficit in this capacity could readily precipitate a decline in model performance.
As illustrated in Figure  5, there is a gradual augmentation in the performance of all models with increasing input length, but as the input extends to considerable lengths (768), some models begin to exhibit diminished performance, which may also be attributed to the amplified noise accompanying the expanded input. In contrast, our model and DLinear consistently demonstrate steady improvements in performance as the input lengthens, exemplifying the superiority of linear models in this context.
5 Conclusions
This paper provides an in-depth exploration of existing solutions for Long-Term Time Series Forecasting (LTSF) tasks within the framework of the Transformer architecture and introduces an innovative approach: Patch-based Decomposed Multi-Layer Perceptron (PDMLP). Through empirical analysis, we demonstrate the efficacy of the patch mechanism in time series forecasting and address the current limitations of Transformer models by designing a novel, simplified architecture. The PDMLP utilizes a simple moving average to separate the smooth components from the noisy residuals in time series, and employs a unique channel mixing strategy to enhance the interchange of semantic information across variables. Additionally, we present the Multi-Scale Patch Embedding method to enable the model to more effectively learn the diverse associations among input sequences. Through extensive experimentation on multiple real-world datasets, PDMLP shows superior performance compared to existing technologies. This research not only validates the potential of reducing complexity and simplifying network structures to improve performance on LTSF tasks but also emphasizes the critical role of cross-variable interactions in enhancing the accuracy of multivariate time series forecasting . We anticipate that the success of PDMLP will not only advance further research in the field of LSTF but also motivate future efforts toward developing models that prioritize efficiency, simplicity, and interpretability. Ultimately, we hope that this research will inspire the creation of more innovative time series forecasting methods focused on addressing specific problems rather than the pursuit of model complexity.
Appendix A Related Work
A.1 Patch-Based Model
Patch-based methods have already demonstrated their immense potential across various data modalities. In the field of Computer Vision (CV), traditional Convolutional Neural Networks (CNNs) naturally segment images into patches and process each patch to learn local features Krizhevsky et al. [2012]. Vision Transformer (ViT) Dosovitskiy et al. [2020] was the first to integrate the concept of patche into the Transformer Vaswani et al. [2017] model, outperforming CNN-based approaches on large datasets and breaking the inductive bias associated with CNNs. Subsequent models such as BEiT Huang et al. [2024] and MAE He et al. [2022] also segment images into patches as input for the Transformer.
In the realm of time series analysis, PatchTST Nie et al. [2022] innovatively segments time series into multiple patches to extract localized semantic information for subsequent Transformer processing. MTSMAE Tang & Zhang [2022] introduced the masked autoencoder method to self-supervised pre-training for time series, learning deep semantic associations through masked patches. iTransformer Liu et al. [2023] further takes a reverse approach by treating an entire time series as a single patch for processing. Additionally, patch methods have been employed within CNN models, such as PatchMixer Gong et al. [2023]. To address the potential loss of temporal information by permutation-invariant self-attention mechanisms, it incorporates a novel permutational variant convolutional structure to preserve temporal information. Relying entirely on depthwise separable convolutions, PatchMixer extracts both local features and global correlations within a single-scale architecture.
Furthermore, TSMixer Ekambaram et al. [2023] draws inspiration from the success of the MLP-Mixer Tolstikhin et al. [2021] model in CV, combining patches with MLP to propose a lightweight, all-MLP neural architecture specifically designed for MTS and representational learning of patch time series. Concurrently, HDMixer aims to capture semantically enriched patches through effective modeling of hierarchical interactions, designing a Length-Extendable Patch adaptable for MTS, enriching the boundary information of patches and alleviating semantic incoherence within sequences.
A.2 Time Series Forecasting
Time series forecasting is integral to many fields and has seen the development of a wide range of models Lim & Zohren [2021]. Classical time series Yule [1927]; Slutzky [1937]; Wold [1938] models remain reliable workhorses for forecasting, with many industrial applications continuing to utilize them. These classic models boast good interpretability and can effectively capture trends and seasonality in data Box & Pierce [1970]. However, their limitations become increasingly apparent as data volume and complexity grow.
Recurrent Neural Networks (RNNs) have also been extensively used in time series forecasting Salinas et al. [2020]. Their recursive structure gives them an innate advantage in handling sequential data, leading deep learning methods to typically rely on RNNs and their variants for developing sequence-to-sequence forecasting paradigms, which have achieved significant advances Rangapuram et al. [2018]; Lim et al. [2020]. CNNs designed for extracting features from images that are principally localized and spatially invariant, have been adapted for time series analysis through the design of multi-layer causal convolutions, ensuring predictions only utilize past information Borovykh et al. [2017]; Lai et al. [2018]; Bai et al. [2018]. Despite meaningful progress with RNNs and their variants, the current algorithms are still incremental sequence-to-sequence prediction models that struggle to forecast long-torm time series accurately Liu et al. [2020]; Qin et al. [2017]; Sutskever et al. [2014].
Transformers have exhibited stellar performance in domains like Natural Language Processing (NLP) Radford et al. [2018]; Devlin et al. [2018]; Radford et al. [2019] and CV Liu et al. [2021b]; He et al. [2016]. In time series tasks, capturing long-torm dependencies and inter-sequence relations is especially critical—a strength of the Transformer architecture Wen et al. [2022]. Most of these models concentrate on innovating new means to simplify the original attention mechanism, thus obtaining enhanced predictive abilities, particularly for lengthier predictions Zhou et al. [2022]; Liu et al. [2021a]; Li et al. [2019]; Kitaev et al. [2020]. While progress has been made with Transformer-based multivariate forecasting models, studies such as Dlinear Zeng et al. [2023] indicate that a simple univariate linear model can surpass complex Transformers. Many subsequent MLP-based models are further exploring the potential of multivariate linear models Li et al. [2023a, b]; Chen et al. [2023]; Ekambaram et al. [2023].
Appendix B Multivariate Time Series
Figure 6: Visualization of Multivariate Time Series Data. We have chosen PCA and t-SNE methods to perform dimensionality reduction on three different types of data. Due to the excessive number of variables in the data, we have opted to visualize only the first 21 dimensions.
(a) The ETTh2 dataset.
(b) The Exchange dataset.
Figure 7: The off-diagonal subplots display the bivariate relationships between multiple variables, while the subplots on the diagonal show the univariate distribution of each variable individually.
To elucidate the interrelationships among variables within multivariate time series data, we employed Principal Component Analysis (PCA) Wold et al. [1987] and T-Distributed Stochastic Neighbor Embedding (T-SNE) Van der Maaten & Hinton [2008] for dimensionality reduction. As illustrated in the Figure  6, it is readily apparent that there is no distinct boundary between data of different variables. Furthermore, on the ETTh2 and Exchange dataset, we selected two distinct variables, assigning one to the x-axis and the other to the y-axis, to more vividly demonstrate the inter-variable associations. It is observable that certain pairs of variables display strong positive or negative correlations (Figure  7(a)). In the Exchange dataset, many variables exhibit a positive correlation with each other (Figure  7(b)).
A natural hypothesis arises: When employing channel mixing methods for forecasting multivariate time series, the interactive effects between different variables can be captured and utilized by the model, which can then leverage these relationships to enhance forecasting accuracy. We posit that the suboptimal performance of previous models employing channel mixing could be attributed to their erroneous usage of cross-variable interaction information. During the time series data collection process, noise can significantly impact the quality of data, resulting in conspicuous noise points (as shown in the Figure  7(a) and  7(b), where some points clearly deviate from the established relationships among variables). Prior channel mixing models based on temporal points were unable to effectively handle such noise, thereby hindering the model’s ability to learn inter-variable correlations.
Appendix C Time Series Decomposition
(a) Calculate the MSE after summing the trend component and the seasonal component.
(b) Sum the MSE calculated separately for the trend component and the seasonal component.
Figure 8: Visualization of input-96-predict-96 results of DLinear on the ETTm1 dataset.
(a) Calculate the MSE after summing the trend component and the seasonal component.
(b) Sum the MSE calculated separately for the trend component and the seasonal component.
Figure 9: Visualization of input-96-predict-96 results of PatchTST on the ETTm1 dataset.
(a) Calculate the MSE after summing the trend component and the seasonal component.
(b) Sum the MSE calculated separately for the trend component and the seasonal component.
Figure 10: Visualization of input-96-predict-192 results of PatchTST on the Exchange dataset.
(a) Calculate the MSE after summing the trend component and the seasonal component.
(b) Sum the MSE calculated separately for the trend component and the seasonal component.
Figure 11: Visualization of input-96-predict-336 results of PatchTST on the Weather dataset.
Time series analysis is a vital approach to understanding complex data. It involves the modeling and forecasting of series characterized by seasonal fluctuations, trends, and other irregular influences. Numerous studies have attempted to deconstruct these composite data sets and extract interpretable structures through decomposition methods, by breaking down time series into their constituent components, to achieve better understanding and prediction Wu et al. [2021]; Zhou et al. [2022]; Tang & Zhang [2023]. We specifically discusses the decomposition method employed in the Autoformer Wu et al. [2021], accompanied by usage strategies of DLinear Zeng et al. [2023] and PatchTST Nie et al. [2022]. The method initially utilizes a moving average kernel to decompose the input raw data into trend and residual (seasonal) components. Two identical yet parameter-independent modules are then applied to each component, with the cumulative features subsequently added to yield a final forecast.
As show in Figure  8(a), it is observable that the trend component predicted by DLinear is comparatively rougher than the actual smooth trend, and its forecast for the seasonal component appears disordered, deviating significantly from the seasonal component of the original data, suggesting minimal correlation. Furthermore, we visualized the trend and seasonal components predicted by the PatchTST model (Figure  9(a)). The trend component forecast by PatchTST appears somewhat smoother than that of DLinear, yet still insufficiently smooth, and it can be stated that the model has not adeptly learned the trend of the original time series.
Table 3: The impact of different components derived from the conventional decomposition method on the predictive performance of the model. Patch is the original method without using decomposition. DLinear and PatchTST_D are the original decomposition method without any changes. DLinear1 and PatchTST1 represent the models where the loss has been modified to calculate the Mean Squared Error (MSE) for the trend and seasonal components separately and then sum them up. DLinear2 and PatchTST2 indicate that the loss is calculated only for the trend component’s MSE, while DLinear3 and PatchTST3 indicate that the loss is calculated only for the seasonal component’s MSE.
Model Prediction Lengths ETTh2 ETTm1 Exchange Weather
MSE MAE MSE MAE MSE MAE MSE MAE
DLinear 96 0.333 0.387 0.345 0.372 0.088 0.218 0.196 0.255
192 0.477 0.476 0.380 0.389 0.176 0.315 0.237 0.296
336 0.594 0.541 0.413 0.413 0.313 0.427 0.283 0.335
720 0.831 0.657 0.474 0.453 0.839 0.695 0.345 0.381
Avg 0.559 0.515 0.403 0.407 0.354 0.414 0.265 0.317
DLinear1 96 0.348 0.397 0.353 0.378 0.086 0.214 0.196 0.256
192 0.469 0.470 0.387 0.395 0.170 0.309 0.238 0.296
336 0.546 0.519 0.417 0.416 0.305 0.417 0.282 0.331
720 0.797 0.641 0.475 0.453 0.831 0.686 0.343 0.383
Avg 0.540 0.507 0.408 0.411 0.348 0.407 0.265 0.317
DLinear2 96 0.406 0.446 0.430 0.432 0.091 0.222 0.213 0.273
192 0.527 0.513 0.459 0.447 0.178 0.319 0.253 0.309
336 0.601 0.556 0.489 0.466 0.306 0.418 0.300 0.349
720 0.859 0.677 0.541 0.498 0.834 0.687 0.359 0.396
Avg 0.598 0.548 0.480 0.461 0.352 0.412 0.281 0.332
DLinear3 96 3.877 1.360 1.403 0.882 4.516 1.563 0.743 0.620
192 3.889 1.346 1.358 0.860 4.395 1.521 0.872 0.670
336 4.055 1.388 1.240 0.823 4.791 1.634 0.828 0.653
720 4.043 1.404 1.334 0.863 4.706 1.622 0.805 0.652
Avg 3.966 1.375 1.334 0.857 4.602 1.585 0.812 0.649
PatchTST 96 0.302 0.348 0.329 0.367 0.088 0.205 0.177 0.218
192 0.388 0.400 0.367 0.385 0.176 0.299 0.225 0.259
336 0.426 0.433 0.399 0.410 0.301 0.397 0.278 0.297
720 0.431 0.446 0.454 0.439 0.901 0.714 0.354 0.348
Avg 0.387 0.407 0.387 0.400 0.367 0.404 0.259 0.281
PatchTST_D 96 0.296 0.345 0.337 0.370 0.089 0.209 0.186 0.224
192 0.373 0.397 0.376 0.386 0.181 0.302 0.232 0.265
336 0.382 0.413 0.397 0.406 0.337 0.421 0.285 0.301
720 0.417 0.443 0.458 0.442 0.919 0.710 0.360 0.351
Avg 0.367 0.400 0.392 0.401 0.382 0.411 0.266 0.285
PatchTST1 96 0.298 0.343 0.339 0.368 0.084 0.203 0.177 0.217
192 0.377 0.390 0.373 0.380 0.177 0.299 0.225 0.260
336 0.389 0.409 0.405 0.402 0.338 0.419 0.282 0.299
720 0.416 0.432 0.466 0.438 0.838 0.687 0.360 0.350
Avg 0.370 0.394 0.396 0.397 0.359 0.402 0.261 0.282
PatchTST2 96 0.365 0.401 0.399 0.410 0.085 0.206 0.192 0.229
192 0.439 0.440 0.437 0.431 0.179 0.302 0.239 0.271
336 0.443 0.452 0.466 0.446 0.330 0.415 0.297 0.311
720 0.471 0.472 0.522 0.477 0.839 0.687 0.373 0.360
Avg 0.430 0.441 0.456 0.441 0.358 0.403 0.275 0.293
PatchTST3 96 0.344 0.370 0.739 0.558 0.148 0.276 0.232 0.280
192 0.425 0.416 0.710 0.549 0.245 0.359 0.282 0.315
336 0.407 0.420 0.739 0.568 0.392 0.459 0.328 0.343
720 0.430 0.439 0.781 0.587 0.944 0.738 0.396 0.384
Avg 0.402 0.411 0.742 0.566 0.432 0.458 0.310 0.331
Additionally, we investigated the influence of traditional time series decomposition methods on the extracted trend and seasonal components concerning model prediction performance. We selected the DLinear and PatchTST models, and chose four datasets including ETTh2, ETTm1, Exchange, and Weather. As shown in Table  3, we can see that models that modify the loss to calculate the Mean Squared Error (MSE) separately for the trend and seasonal components and then sum them achieve better performance. This indicates that computing the loss separately constrains the model to learn these components more in line with our expectations. However, for PatchTST, it seems that decomposing the trend and seasonal components with traditional decomposition methods and then calculating the loss after summing them does not appear to bring much performance improvement. Moreover, as can be seen in Figure  9, the model does not seem to have learned much information, especially with components like the trend which appear to be very smooth. This suggests that traditional methods of time series decomposition may not be very feasible. In some datasets (e.g., Figure  10), the model did not learn any information at all—both trend and seasonality could essentially be described as predicting a constant.
Appendix D Attention score map of Transformer
(a) ECL
(b) ETTm1
(c) Traffic
Figure 12: The self-attention scores from a 2-layer Transformer trained on different datasets, showing that time series data tends to be patch.
The traditional Transformer does not adopt the idea of patche but instead directly embeds the time series at the point-level. Much research indicates that this approach tends to lead to the Transformer processing a large amount of redundant features Tang & Zhang [2022], resulting in lower attention values for certain time points, exhibiting a long-tail distribution with attention scores clearly showing a form of blockiness Zhou et al. [2021]. Figure  12 shows a typical attention score map of the original Transformer for time series forecasting. We selected three datasets of different sizes based on different sampling scales and trained a two-layer Transformer with eight heads, ultimately selecting the attention scores from the second layer. We can see that attention values tend to patch, i.e., close data points have similar attention weights. This suggests that directly computing attention on time points is not advisable for Transformers. However, as shown in Figure  1, creating patches from the input time series can only mitigate this blockiness and does not completely resolve it.
Appendix E Experiments
We further delved into the efficacy of each component of PDMLP and its variants when applied to the LSTF task, and investigated the sensitivity of PDMLP to a range of hyperparameters.
E.1 Datasets
Table 4: Detailed dataset descriptions. $Dim$ denotes the number of variables in each dataset. $Dataset$ $Size$ indicates the total number of time points split across the training, validation, and test sets, respectively. $Prediction$ $Length$ refers to the future time points to be forecasted, with each dataset encompassing four forecasting settings. $Frequency$ denotes the sampling interval of time points for this dataset.
Dataset Dim Prediction Length Dataset Size Frequency Information
ETTh1, ETTh2 7 {96, 192, 336, 720} (8545, 2881, 2881) Hourly Electricity
ETTm1, ETTm2 7 {96, 192, 336, 720} (34465, 11521, 11521) 15min Electricity
Exchange 8 {96, 192, 336, 720} (5120, 665, 1422) Daily Economy
Weather 21 {96, 192, 336, 720} (36792, 5271, 10540) 10min Weather
ECL 321 {96, 192, 336, 720} (18317, 2633, 5261) Hourly Electricity
Traffic 862 {96, 192, 336, 720} (12185, 1757, 3509) Hourly Transportation
Solar-Energy 137 {96, 192, 336, 720} (36601, 5161, 10417) 10min Energy
We evaluated the performance of our proposed PDMLP on 8 commonly used LSTF benchmark datasets: Solar Energy Lai et al. [2018], Weather, Traffic, Electricity (ECL) Wu et al. [2021], and 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2) Zhou et al. [2021]. The details of datasets are provided in Table  4.
  • ETT captured data incorporating seven factors pertinent to electric power transformers from July 2016 to July 2018. It comprises four subsets: ETTh1 and ETTh2, which record data hourly, and ETTm1 and ETTm2, which record data every 15 minutes.
  • Exchange compiled panel data of daily exchange rates from eight countries spanning the period from 1990 to 2016.
  • Weather encompasses twenty-one meteorological factors collected every 10 minutes from the Max Planck Institute for Biogeochemistry’s weather station in the year 2020.
  • ECL logged hourly electricity consumption data for 321 customers.
  • Traffic gathered data on hourly roadway occupancy rates measured by 862 sensors on the highways of the San Francisco Bay Area from January 2015 to December 2016.
  • Solar Energy chronicled the power output from 137 photovoltaic power stations in the year 2006, with samples taken every 10 minutes.
E.2 Implementation Details
Table 5: Default setting.
config value
optimizer Adam
learning rate $10^{-4}$
batch size 32
total epoch 20
early stopping 5
input length 96
MLP layer 1
d_model 1024
patch size 48,24,12,6
dropout 0.1
loss function MSE
All experiments were implemented using PyTorch Paszke et al. [2019] and executed on a single NVIDIA V100 16GB GPU. Table  5 presents the experimental setup. Optimization of the model was carried out using the Aadm optimizer Kingma & Ba [2014], with an initial learning rate of $10^{-4}$ and L2 loss. We standardized the batch size to 32, fixed the number of training epochs to 20, and employed an early stopping criterion, where if the loss on the validation set did not decrease for five consecutive epochs, training would be terminated prematurely. In our proposed model, the number of MLP layers, denoted as N, was set to 1. The dimension of series representations d_model was established at 1024. Unless otherwise specified, the input sequence length for the models was set to 96. For PDMLP, a consistent dropout probability of 0.1 was applied across all positions. Regarding Multi-Scale Patch Embedding, we selected four patch sizes: [48, 24, 12, 6].
All baseline models to which we compared our approach were based on the iTransformer Liu et al. [2023] and TimesNet Wu et al. [2022] repositories, constructed atop configurations provided by the original papers or official codes of each respective model. Furthermore, except for results derived from the iTransformer repository, all other results were obtained by conducting six separate runs with different random seeds, and the mean of these six outcomes was reported.
E.3 Hyperparameter Sensitivity
Figure 13: Hyperparameter sensitivity with respect to the learning rate, the number of MLP blocks, and the hidden dimension of Embedding tokens. The results are recorded with the lookback window length $L=96$ and the forecast window length $T=\{96,192\}$
We evaluated the hypersensitivity of the PDMLP concerning the following parameters: learning rate (lr), the number of blocks (layers) N in the MLP, and the hidden dimension D of the embedding. The results are illustrated in Figure  13. We observed that the performance of PDMLP is not particularly sensitive to these hyperparameters, as evidenced by the relatively inconspicuous variations in performance. It is advisable that both the number of blocks and the hidden dimension size not be excessively large, which may be attributable to the increase in the total number of parameters that require learning with larger values of these hyperparameters. However, the hidden dimension size should not be excessively small either, to avoid a disproportionate compression of features into an insufficiently capacious latent vector.
E.4 Layer Normalization and Batch Normalization
Table 6: The Impact of Different Normalization Techniques on the Predictive Performance of Various Models. The notations 1 and 2 are employed to signify that Layer Normalization (LN) and Batch Normalization (BN) are used in the models, respectively. In the case of PDMLP, which incorporates two successive normalization layers, PDMLP3 refers to a configuration wherein the first normalization layer is composed of BN and the second of LN; conversely, PDMLP4 indicates that the initial normalization layer utilizes LN followed by a BN layer. We compare the performance of identical models employing disparate normalization techniques, with the optimal outcomes highlighted in bold typeface.
Dateset ETTh2 ETTm1 Exchange
Model Metric 96 192 336 720 Avg 96 192 336 720 Avg 96 192 336 720 Avg
PDMLP1 MSE 0.288 0.371 0.412 0.421 0.373 0.316 0.357 0.396 0.462 0.382 0.085 0.174 0.317 0.820 0.349
MAE 0.340 0.396 0.429 0.441 0.401 0.357 0.378 0.402 0.439 0.394 0.204 0.298 0.411 0.681 0.398
PDMLP2 MSE 0.292 0.379 0.431 0.436 0.385 0.319 0.365 0.402 0.472 0.390 0.088 0.178 0.326 0.862 0.364
MAE 0.344 0.400 0.438 0.450 0.408 0.358 0.383 0.406 0.444 0.398 0.209 0.302 0.413 0.699 0.406
PDMLP3 MSE 0.300 0.386 0.426 0.434 0.387 0.321 0.365 0.407 0.472 0.391 0.089 0.179 0.326 0.855 0.362
MAE 0.347 0.398 0.435 0.447 0.407 0.360 0.382 0.407 0.443 0.398 0.207 0.301 0.413 0.698 0.405
PDMLP4 MSE 0.306 0.389 0.428 0.438 0.390 0.323 0.371 0.404 0.472 0.393 0.088 0.182 0.338 0.873 0.370
MAE 0.354 0.404 0.434 0.452 0.411 0.362 0.387 0.409 0.445 0.401 0.206 0.303 0.423 0.703 0.409
iTransformer1 MSE 0.297 0.380 0.428 0.427 0.383 0.334 0.377 0.426 0.491 0.407 0.086 0.177 0.331 0.847 0.360
MAE 0.349 0.400 0.432 0.445 0.407 0.368 0.391 0.420 0.459 0.410 0.206 0.299 0.417 0.691 0.403
iTransformer2 MSE 0.318 0.379 0.427 0.438 0.391 0.336 0.392 0.445 0.506 0.420 0.093 0.186 0.336 0.886 0.375
MAE 0.350 0.398 0.433 0.448 0.407 0.371 0.399 0.429 0.466 0.416 0.215 0.310 0.420 0.710 0.414
PatchTST1 MSE 0.299 0.372 0.415 0.426 0.378 0.335 0.374 0.403 0.462 0.394 0.086 0.180 0.332 0.826 0.356
MAE 0.347 0.397 0.430 0.444 0.405 0.370 0.390 0.409 0.447 0.404 0.208 0.303 0.417 0.680 0.402
PatchTST2 MSE 0.302 0.388 0.426 0.431 0.387 0.329 0.367 0.399 0.454 0.387 0.088 0.176 0.301 0.901 0.367
MAE 0.348 0.400 0.433 0.446 0.407 0.367 0.385 0.410 0.439 0.400 0.205 0.299 0.397 0.714 0.404
TimesNet1 MSE 0.340 0.402 0.452 0.462 0.414 0.338 0.374 0.410 0.478 0.400 0.107 0.226 0.367 0.964 0.416
MAE 0.374 0.414 0.452 0.468 0.427 0.375 0.387 0.411 0.450 0.406 0.234 0.344 0.448 0.746 0.443
TimesNet2 MSE 0.335 0.432 0.470 0.515 0.438 0.393 0.500 0.553 0.635 0.520 0.108 0.247 0.515 1.301 0.543
MAE 0.372 0.426 0.466 0.496 0.440 0.413 0.456 0.486 0.527 0.471 0.236 0.361 0.529 0.877 0.501
Normalization is a critical processing technique to reduce internal covariate shift while training deep neural networks. Although the choice between Layer Normalization (LN) Ba et al. [2016] and Batch Normalization (BN) Ioffe & Szegedy [2015] depends on the specific task, and PatchTST has demonstrated certain advantages of BN in time series datasets, LN generally shows better robustness and compatibility in sequence tasks.
We selected four models, PDMLP, iTransformer, PatchTST, and TimesNet, to contrast their performance differences with the application of LN and BN over three datasets: ETTh2, ETTm1, and Exchange. Among these, as stated in its original publication, PatchTST emphasizes the use of BN. However, the TimesNet approach, which is predicated on CNN architectures, primarily utilizes LN. Table  6 illustrates that the BN method did not achieve the surprisingly high performance delineated in the PatchTST paper. On the contrary, LN proved to yield superior results in the majority of the experiments conducted. This suggests that, relative to BN, LN remains more adept at handling time-series tasks. For many models, the performance difference between LN and BN is not particularly significant, but for TimesNet, BN performs 30% lower than LN.
E.5 Variants of PDMLP
Figure 14: A variant of Intra-Variable MLP or Inter-Variable MLP involves modifying a single fully-connected layer into two fully-connected layers. d_model and d_FN are the dimensions of the latent vector
In reference to the Intra-Variable MLP and the Inter-Variable MLP, our previous work employed a single fully connected layer; however, we have now flexibly adapted them as dual-layer fully connected structures, as depicted in Figure  14. Moreover, in the previous versions of PDMLP, we utilized the same MLP Layer to process both the Smooth and Residual components independently. Currently, we have flexibly implemented variant MLP Layer to address these disparate components. As illustrated in Table  7, the $\ast$ denotes the previous single-layer MLP, $\Diamond$ denotes the dual-layer MLP with feature number $d\_FN=d\_model/\times 2$, and $\triangle$ indicates a dual-layer MLP with $d\_FN=d\_model/2$. Re-order signifies a change in the sequence of Intra-Variable MLP and Inter-Variable MLP operations, prompting the model to initiate cross-variable interactions prior to the within-variable interactions.
Upon examination, it is apparent that different PDMLP variants do not significantly alter the performance outcome. However, the integration of the dual-layer MLP leads to an increase in the number of learning parameters within the model. Even though the $\triangle$ type dual-layer MLP does not modify the quantity of parameters, it potentially can result in information loss due to dimensionality compression. Additionally, it is noteworthy that altering the sequence of variable interactions, as seen in the $Re-order$ configuration, yields inferior performance. This suggests that intra-variable interactions following the cross-variable interactions may lead to a loss of the information obtained from the cross-variable interactions.
Table 7: Different variants of PDMLP. The part before the $+$ indicates the type of MLP used for Intra-Variable, while the part after the $+$ indicates the type of MLP adopted for Inter-Variable.
Smooth MLP Layer Residual MLP Layer Prediction Lengths ETTh2 ETTm1 Exchange Weather
MSE MAE MSE MAE MSE MAE MSE MAE
$\ast$+$\ast$ + ∗ $\ast$+$\ast$ + ∗ 96 0.288 0.340 0.316 0.357 0.085 0.204 0.157 0.209
192 0.371 0.396 0.357 0.378 0.174 0.298 0.206 0.251
336 0.412 0.429 0.396 0.402 0.317 0.411 0.267 0.293
720 0.421 0.441 0.462 0.439 0.820 0.681 0.347 0.346
Avg 0.373 0.401 0.382 0.394 0.349 0.398 0.244 0.274
Re-order Re-order 96 0.315 0.361 0.320 0.359 0.093 0.216 0.165 0.212
192 0.388 0.407 0.368 0.385 0.192 0.313 0.216 0.256
336 0.436 0.442 0.407 0.410 0.357 0.433 0.270 0.295
720 0.460 0.462 0.486 0.451 0.955 0.732 0.351 0.347
Avg 0.400 0.418 0.395 0.401 0.399 0.424 0.251 0.278
$\ast$+$\Diamond$ + ◇ $\ast$+$\Diamond$ + ◇ 96 0.300 0.349 0.320 0.360 0.096 0.209 0.161 0.208
192 0.391 0.404 0.367 0.385 0.189 0.309 0.211 0.253
336 0.428 0.434 0.400 0.407 0.340 0.421 0.267 0.294
720 0.438 0.452 0.481 0.448 0.907 0.715 0.349 0.347
Avg 0.389 0.410 0.392 0.400 0.383 0.414 0.247 0.276
$\Diamond$+$\ast$ + ∗ $\Diamond$+$\ast$ + ∗ 96 0.310 0.356 0.326 0.364 0.090 0.211 0.163 0.210
192 0.393 0.407 0.374 0.390 0.186 0.308 0.211 0.252
336 0.435 0.438 0.410 0.412 0.342 0.424 0.267 0.293
720 0.441 0.452 0.486 0.452 0.897 0.711 0.347 0.346
Avg 0.395 0.413 0.399 0.405 0.379 0.414 0.247 0.275
$\Diamond$+$\Diamond$ + ◇ $\Diamond$+$\Diamond$ + ◇ 96 0.303 0.350 0.321 0.360 0.091 0.210 0.161 0.208
192 0.393 0.405 0.372 0.387 0.185 0.306 0.210 0.252
336 0.430 0.439 0.407 0.411 0.348 0.427 0.268 0.394
720 0.438 0.451 0.482 0.450 0.927 0.719 0.347 0.346
Avg 0.391 0.411 0.396 0.402 0.388 0.416 0.247 0.300
$\ast$+$\Diamond$ + ◇ $\ast$+$\triangle$ + △ 96 0.297 0.347 0.325 0.362 0.090 0.210 0.163 0.210
192 0.386 0.400 0.365 0.382 0.185 0.306 0.212 0.253
336 0.437 0.438 0.395 0.403 0.345 0.424 0.269 0.294
720 0.439 0.450 0.472 0.445 0.904 0.715 0.348 0.346
Avg 0.390 0.409 0.389 0.398 0.381 0.414 0.248 0.276
$\ast$+$\Diamond$ + ◇ $\ast$+$\ast$ + ∗ 96 0.303 0.351 0.325 0.362 0.091 0.210 0.164 0.210
192 0.387 0.402 0.367 0.384 0.185 0.306 0.213 0.253
336 0.428 0.435 0.402 0.407 0.345 0.425 0.268 0.294
720 0.445 0.454 0.476 0.447 0.909 0.715 0.347 0.345
Avg 0.391 0.411 0.393 0.400 0.383 0.414 0.248 0.276
$\ast$+$\triangle$ + △ $\ast$+$\triangle$ + △ 96 0.301 0.349 0.322 0.361 0.089 0.207 0.167 0.213
192 0.387 0.399 0.363 0.382 0.186 0.306 0.213 0.253
336 0.426 0.434 0.402 0.408 0.342 0.422 0.271 0.395
720 0.440 0.453 0.473 0.446 0.873 0.702 0.348 0.345
Avg 0.389 0.409 0.390 0.399 0.373 0.409 0.250 0.302
$\ast$+$\ast$ + ∗ $\ast$+$\triangle$ + △ 96 0.298 0.348 0.322 0.360 0.091 0.211 0.166 0.213
192 0.390 0.403 0.366 0.384 0.185 0.307 0.214 0.254
336 0.426 0.434 0.397 0.405 0.343 0.424 0.271 0.295
720 0.441 0.452 0.475 0.446 0.916 0.716 0.348 0.346
Avg 0.389 0.409 0.390 0.399 0.384 0.415 0.250 0.277
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
[…]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [No]
[…]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
[…]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
[…]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
[…]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
[…]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
[…]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
[…]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
[…]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
[…]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
[…]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
[…]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
[…]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
[…]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
[…]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.